OpenStudy (anonymous):

A rectangular storage container with an open top is to have a volume of 26 cubic meters. The length of its base is twice the width. Material for the base costs 13 dollars per square meter. Material for the sides costs 7 dollars per square meter. Find the cost of materials for the cheapest such container.

OpenStudy (cwrw238):

area of base = l*w = 2w*w = 2w^2 cost of base = 13*2w^2 = 26w^2 area of sides = 2*2wh + 2*wh = 6wh cost of the sides = 7*6wh = 42wh total cost C = 26w^2 + 42wh

OpenStudy (cwrw238):

volume 26 = 2w^2h h = 26/2w^2 = 13/w^2 substitute for h into the formula for C: C = 26w^2 + 42w * 13/w^2 C = 26w^2 + 546/w

OpenStudy (cwrw238):

we need to find the minimum value of this function

OpenStudy (cwrw238):

now differentiate C with respect to w and equate to zero solve for w then plug in this value for w into the formula for C and you have it.